OpenStudy (anonymous):

Mr. Bird’s son is on a coed soccer team with Mrs. Twiddy’s daughters for a total of 36 team members. The number of girls is 3 less than double the number of boys. How many of the players are boys and how many are girls? Write a system of equations to model the problem. Solve the system by substitution, show your work. Give your final answer in a complete sentence. @perl

OpenStudy (perl):

Let `g` stand for number of girls, let `b` stand for number of boys.

OpenStudy (perl):

B + G = 36 G = 2B - 3

OpenStudy (anonymous):

ok

OpenStudy (perl):

can you try to solve that now (use substitution)

OpenStudy (anonymous):

i dont know how

OpenStudy (perl):

we know that G = 2B - 3 , so we can substitute that in the first equation

OpenStudy (anonymous):

ok how do u do that lol

OpenStudy (perl):

B + G = 36 G = 2B - 3 therefore B + (2B -3 ) = 36

OpenStudy (anonymous):

ok

OpenStudy (perl):

you can use the equation $$ \Large { n_1 \sin\theta_1 = n_2 \sin \theta_2 } $$

OpenStudy (perl):

oops wrong post

OpenStudy (anonymous):

lol

OpenStudy (perl):

does that make sense, how i got B + (2B -3 ) = 36

OpenStudy (anonymous):

yes

OpenStudy (perl):

ok collect like terms and solve that

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

3B-3+36?

OpenStudy (perl):

correct

OpenStudy (perl):

3B - 3 = 36 now add 3 to both sides

OpenStudy (anonymous):

\[3B - 3 = 36 \]

OpenStudy (anonymous):

3=33?

OpenStudy (anonymous):

@perl

OpenStudy (perl):

$$ \Large{ \\~\\3B - 3 = 36 \\~\\\\ 3B = 39 \\~\\\\ B = \frac {39}{3} }$$